# VertiBench: Advancing Feature Distribution Diversity in Vertical Federated Learning Benchmarks

### v1

@@ 7 Conclusion
In this study, we introduce VertiBench, a versatile benchmarking framework for Vertical Federated Learning (VFL). VertiBench facilitates the synthetic generation of diverse VFL datasets from a single global set, thereby enabling a comprehensive performance assessment of VFL algorithms across a wide spectrum of application domains. Our empirical results reveal potential significant variations in algorithm performance under different data partition scenarios, underscoring the importance of our benchmark. Additionally, we contribute a new real-world VFL dataset, addressing a deficit in image-image VFL datasets. This study highlights the necessity of examining VFL algorithms under diverse data distribution conditions, providing a crucial trajectory for future research.
@@ Appendix A Proof
$\log\mathcal{P}(y|X_{K},...,X_{1})=\sum_{i=1}^{K}\log{\frac{\mathcal{P}(y|X_{k},...,X_{1})}{\mathcal{P}(y|X_{k-1},...,X_{1})}}+\log\mathcal{P}(y)$ (4)
@@
$\;\mathcal{P}(y|X_{K},...,X_{1})$ (5)
$=$ $\;\frac{\mathcal{P}(y,X_{K},...,X_{1})}{\mathcal{P}(X_{K},...,X_{1})}$
$=$ $\;\frac{\mathcal{P}(y)\mathcal{P}(X_{1}|y)\prod_{k=2}^{K}{\mathcal{P}(X_{k}|y,X_{k-1},...,X_{1})}}{\mathcal{P}(X_{1})\prod_{k=2}^{K}{\mathcal{P}(X_{k}|X_{k-1},...,X_{1})}}$
$=$ $\;\mathcal{P}(y)\frac{\mathcal{P}(X_{1}|y)}{\mathcal{P}(X_{1})}\prod_{k=2}^{K}{\frac{\mathcal{P}(X_{k}|y,X_{k-1},...,X_{1})}{\mathcal{P}(X_{k}|X_{k-1},...,X_{1})}}$
@@
$c_{k}=\log\frac{\mathcal{P}(X_{k}|y,X_{k-1},...,X_{1})}{\mathcal{P}(X_{k}|X_{k-1},...,X_{1})},\quad c_{1}=\log\frac{\mathcal{P}(X_{1}|y)}{\mathcal{P}(X_{1})}$ (6)
@@
$c_{k}$ $=\;\frac{\mathcal{P}(X_{k}|y,X_{k-1},...,X_{1})}{\mathcal{P}(X_{k}|X_{k-1},...,X_{1})}$ (8)
$=\;\frac{\mathcal{P}(X_{k},y|X_{k-1},...,X_{1})}{\mathcal{P}(X_{k}|X_{k-1},...,X_{1})\mathcal{P}(y|X_{k-1},...,X)}$
$=\;\frac{\mathcal{P}(y|X_{k},...,X_{1})}{\mathcal{P}(y|X_{k-1},...,X_{1})}$
@@
$\log\mathcal{P}(y|X_{K},...,X_{1})=\sum_{i=1}^{K}\log{\frac{\mathcal{P}(y|X_{k},...,X_{1})}{\mathcal{P}(y|X_{k-1},...,X_{1})}}+\log\mathcal{P}(y)$ (9)
@@
Consider a feature index set $\mathcal{A}=\{1,2,...,m\}$ and a characteristic function $v:2^{\mathcal{A}}\rightarrow\mathbb{R}$ such that $v(\emptyset)=0$. Let $\phi_{j}(v)$ denote the importance of the $j$-th feature on $v$ such that $\sum_{j=1}^{m}\phi(j)=v(\mathcal{A})$. Assume that the indices in $\mathcal{A}$ are randomly distributed to $K$ parties with probabilities $r_{1},...,r_{K}$ where $\sum_{i=1}^{K}{r_{i}}=1$. Given budgets $b_{i}=r_{i}v(\mathcal{A})$, let $Z_{i}$ be the sum of feature importance for party $i$. Then, we have $\forall i\in[1,K],\mathbf{E}[Z_{i}]=b_{i}$ and $\mathbf{E}[Z_{i}]\propto r_{i}$.
@@
$\mathbf{E}[Z_{i}]=\sum_{j=1}^{m}\mathbf{E}[Y_{ij}]=\sum_{j=1}^{m}\phi_{j}(v)r_{i}=r_{i}\sum_{j=1}^{m}\phi(j)$ (11)
@@
$\mathbf{E}[Z_{i}]=b_{i}$ (12)
@@
Suppose we have variables $X_{1},\dots,X_{K}$ following the Dirichlet distribution, denoted as $\text{Dir}(\overline{\alpha},\dots,\overline{\alpha})$. Leveraging the inherent properties of the Dirichlet distribution, we can formulate the variance $\text{Var}(X_{i})$ for all $i\in[1,K]$ as
$\text{Var}(X_{i})=\frac{K-1}{K^{2}(K\overline{\alpha}+1)}$ (14)
@@
$\mathbb{E}[\text{Var}(X_{i})]=\frac{K-1}{K^{2}(K\overline{\alpha}+1)}=\sigma$ (15)
Recognizing that for a Dirichlet distribution $\sigma>0$ holds, we can transform the above equation to express $\overline{\alpha}$ in terms of $\sigma$:
$\overline{\alpha}=\frac{K-1-K^{2}\sigma}{K^{3}\sigma}$ (16)
@@ Appendix B Details of VFL Algorithms
In this section, we provide a detailed comparison of the existing VFL algorithms as an extension of Table 1.

### v2

@@ 5 Conclusion
We introduce VertiBench, a refined benchmarking tool for Vertical Federated Learning (VFL), adept at generating a variety of synthetic VFL datasets from a single global dataset. The scope of VertiBench extends beyond the confines of existing uniform and real scopes, shedding light on VFL scenarios previously unexplored. Our findings underscore performance variations under diverse data partitions, emphasizing the need to evaluate VFL algorithms across varied feature splits for enhanced insights into their real-world applicability.
6 Reproducibility Statement
The code for this study is accessible via an anonymous repository Anonymized (2023a), accompanied by a README.md file that provides guidelines for environment setup and result reproduction. Comprehensive proofs of all theoretical results are meticulously detailed in Appendix A. Further, Appendix F offers a detailed description of dataset specifications and hyperparameter configurations.
@@ Appendix A Proof
$\log\mathbb{P}(y|X_{K},...,X_{1})=\sum_{i=1}^{K}\log{\frac{\mathbb{P}(y|X_{k},...,X_{1})}{\mathbb{P}(y|X_{k-1},...,X_{1})}}+\log\mathbb{P}(y)$ (4)
@@
$\;\mathbb{P}(y|X_{K},...,X_{1})$ (5)
$=$ $\;\frac{\mathbb{P}(y,X_{K},...,X_{1})}{\mathbb{P}(X_{K},...,X_{1})}$
$=$ $\;\frac{\mathbb{P}(y)\mathbb{P}(X_{1}|y)\prod_{k=2}^{K}{\mathbb{P}(X_{k}|y,X_{k-1},...,X_{1})}}{\mathbb{P}(X_{1})\prod_{k=2}^{K}{\mathbb{P}(X_{k}|X_{k-1},...,X_{1})}}$
$=$ $\;\mathbb{P}(y)\frac{\mathbb{P}(X_{1}|y)}{\mathbb{P}(X_{1})}\prod_{k=2}^{K}{\frac{\mathbb{P}(X_{k}|y,X_{k-1},...,X_{1})}{\mathbb{P}(X_{k}|X_{k-1},...,X_{1})}}$
@@
$c_{k}=\log\frac{\mathbb{P}(X_{k}|y,X_{k-1},...,X_{1})}{\mathbb{P}(X_{k}|X_{k-1},...,X_{1})},\quad c_{1}=\log\frac{\mathbb{P}(X_{1}|y)}{\mathbb{P}(X_{1})}$ (6)
@@
$c_{k}$ $=\;\frac{\mathbb{P}(X_{k}|y,X_{k-1},...,X_{1})}{\mathbb{P}(X_{k}|X_{k-1},...,X_{1})}$ (8)
$=\;\frac{\mathbb{P}(X_{k},y|X_{k-1},...,X_{1})}{\mathbb{P}(X_{k}|X_{k-1},...,X_{1})\mathbb{P}(y|X_{k-1},...,X)}$
$=\;\frac{\mathbb{P}(y|X_{k},...,X_{1})}{\mathbb{P}(y|X_{k-1},...,X_{1})}$
@@
$\log\mathcal{P}(y|X_{K},...,X_{1})=\sum_{i=1}^{K}\log{\frac{\mathbb{P}(y|X_{k},...,X_{1})}{\mathbb{P}(y|X_{k-1},...,X_{1})}}+\log\mathbb{P}(y)$ (9)
Proposition 2.
[…]
Proof.
In the context of the correlation matrix of $\mathbf{X}$, Pcor measures the standard deviation of singular values, while mcor Taylor [2020] measures the standard deviation of eigenvalues. Given that the correlation matrix is symmetric positive semi-definite, each eigenvalue corresponds to a singular value, resulting in Pcor equating to mcor. ∎
Proposition 3.
[…]
Proof.
Denote the correlation matrix between $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ as $C$, the singular values of $C$ as $\Sigma=[\sigma_{1},\dots,\sigma_{d}]$. Then
$\text{Pcor}(\mathbf{X}_{1},\mathbf{X}_{2})=\sqrt{\frac{1}{d}\text{Var}(\Sigma)}=\sqrt{\frac{1}{d}(\mathbb{E}(\Sigma^{2})-\mathbb{E}^{2}(\Sigma))}$ (10)
Thus,
$0\leq\text{Pcor}(\mathbf{X}_{1},\mathbf{X}_{2})\leq\sqrt{\frac{1}{d}\mathbb{E}(\Sigma^{2})}$ (11)
Given that every element in the correlation matrix is in range $[-1,1]$
$\mathbb{E}(\Sigma^{2})=\text{tr}(C^{T}C)/d\leq d^{2}/d\leq d$ (12)
Thus, $\text{Pcor}(\mathbf{X}_{1},\mathbf{X}_{2})\in[0,1]$
@@
[…]
The lower bound occurs if and only if $\mathbf{X}_{1}$ comprises all features of either $\mathbf{D}_{1}$ or $\mathbf{D}_{2}$, with $\mathbf{X}_{2}$ containing the remaining features. The upper bound occurs if and only if $\mathbf{X}_{1}$ holds $m$ features from both $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$, with $\mathbf{X}_{2}$ holding the remaining $m$ features from $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$.
Proof.
Consider two sets of features, $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, derived from datasets $\mathbf{D}_{1}$ and $\mathbf{D}_{2}$. Let $\mathbf{X}_{1}$ contain $u$ features from $\mathbf{D}_{1}$ and $v$ features from $\mathbf{D}_{2}$, where $u,v\in[0,m]$ and $u+v=m$. The set $\mathbf{X}_{2}$ comprises the remaining features, specifically $v$ features from $\mathbf{D}_{1}$ and $u$ features from $\mathbf{D}_{2}$. The inner-party correlation matrices $\mathbf{C}_{11}$ for $\mathbf{X}_{1}$ and $\mathbf{C}_{22}$ for $\mathbf{X}_{2}$ are given by:
$\mathbf{C}_{11}=\begin{bmatrix}\mathbb{1}^{u\times u}&\mathbb{0}^{u\times v}\\ \mathbb{0}^{v\times u}&\mathbb{1}^{v\times v}\end{bmatrix},\mathbf{C}_{22}=\begin{bmatrix}\mathbb{1}^{v\times v}&\mathbb{0}^{v\times u}\\ \mathbb{0}^{u\times v}&\mathbb{1}^{u\times u}\end{bmatrix}$ (13)
To calculate the $\text{Pcor}(\mathbf{X}_{1},\mathbf{X}_{1})$ and $\text{Pcor}(\mathbf{X}_{2},\mathbf{X}_{2})$, we need to determine the singular values of $\mathbf{C}_{11}$ and $\mathbf{C}_{22}$. As these are symmetric matrices, their singular values are the absolute values of their eigenvalues. We will illustrate the process using $\mathbf{C}_{11}$, with $\mathbf{C}_{22}$ following a similar pattern.
The characteristic equation of $\mathbf{C}_{11}$ is
$\begin{vmatrix}\mathbf{U}_{11}^{u\times u}&\mathbb{0}^{u\times v}\\ \mathbb{0}^{v\times u}&\mathbf{V}_{11}^{v\times v}\end{vmatrix}=0$ (14)
where $\mathbf{U}_{11}^{u\times u}$ and $\mathbf{V}_{11}^{v\times v}$ are matrices defined as
$\mathbf{U}_{11}^{u\times u}=\mathbf{V}_{11}^{v\times v}=\begin{bmatrix}1-\lambda&1&\cdots&1\\ 1&1-\lambda&\cdots&1\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1-\lambda\end{bmatrix}$
Equation 14 simplifies to $\text{det}(\mathbf{U}_{11}^{u\times u})\text{det}(\mathbf{V}_{11}^{v\times v})=0$. The roots of $\text{det}(\mathbf{U}_{11}^{u\times u})=0$ and $\text{det}(\mathbf{V}_{11}^{v\times v})=0$ are the eigenvalues of $\mathbb{1}^{u\times u}$ and $\mathbb{1}^{v\times v}$ respectively. Since $\mathbb{1}^{u\times u}$ and $\mathbb{1}^{v\times v}$ are rank-1 matrices, they each have only one non-zero eigenvalue, $u$ and $v$ respectively. Therefore, $\mathbf{C}_{11}$ has two non-zero eigenvalues: $u$ and $v$. The Pcor for $\mathbf{X}_{1}$ is
$\text{Pcor}(\mathbf{X}_{1},\mathbf{X}_{1})=\frac{1}{\sqrt{m}}\sqrt{\frac{u^{2}+v^{2}}{m-1}}$ (15)
A similar calculation for $\mathbf{C}_{22}$ yields:
$\text{Pcor}(\mathbf{X}_{2},\mathbf{X}_{2})=\frac{1}{\sqrt{m}}\sqrt{\frac{u^{2}+v^{2}}{m-1}}$ (16)
Next, we consider the inter-party correlation matrices $\mathbf{C}_{12}$ and $\mathbf{C}_{21}$
$\mathbf{C}_{12}=\begin{bmatrix}\mathbb{1}^{u\times v}&\mathbb{0}^{u\times u}\\ \mathbb{0}^{v\times v}&\mathbb{1}^{v\times u}\end{bmatrix},\mathbf{C}_{21}=\mathbf{C}_{12}^{T}=\begin{bmatrix}\mathbb{1}^{v\times u}&\mathbb{0}^{v\times v}\\ \mathbb{0}^{u\times u}&\mathbb{1}^{u\times v}\end{bmatrix}$ (17)
Since $\mathbf{C}_{21}=\mathbf{C}_{12}^{T}$, they share the same singular values. A key principle in linear algebra states that the singular values of a matrix remain invariant under rotation. This invariance arises because rotating a matrix is mathematically equivalent to multiplying it by an orthogonal matrix, a process that does not alter its singular values. By rotating $\mathbf{C}_{12}$ and $\mathbf{C}_{21}$ 90 degrees counter-clockwise, we obtain symmetric matrices $\mathbf{C}_{12}^{\prime}$ and $\mathbf{C}_{21}^{\prime}$, whose singular values are the absolute values of their eigenvalues. As $\mathbf{C}_{12}^{\prime}$ is a rank-2 matrix, it has two non-zero eigenvalues.
The characteristic equation of $\mathbf{C}_{12}^{\prime}$ is
$\begin{vmatrix}\mathbf{U}_{12}^{u\times u}&\mathbb{1}^{u\times v}\\ \mathbb{1}^{v\times u}&\mathbf{V}_{12}^{v\times v}\end{vmatrix}=0$ (18)
where $\mathbf{U}_{12}^{u\times u}$ and $\mathbf{V}_{12}^{v\times v}$ are defined as
$\mathbf{U}_{12}^{u\times u}=\begin{bmatrix}-\lambda&0&\cdots&0\\ 0&-\lambda&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-\lambda\end{bmatrix}^{u\times u},\mathbf{V}_{12}^{v\times v}=\begin{bmatrix}-\lambda&0&\cdots&0\\ 0&-\lambda&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-\lambda\end{bmatrix}^{v\times v}$ (19)
Since we focused on non-zero eigenvalues ($\lambda\neq 0$), $\mathbf{U}_{12}^{u\times u}$ is invertible. According to the property of block matrices, Eq. 18 is equivalent to
$\det(\mathbf{U}_{12}^{u\times u})\det(\mathbf{V}_{12}^{v\times v}-\mathbb{1}^{v\times u}({\mathbf{U}_{12}^{-1}})^{u\times u}\mathbb{1}^{u\times v})$ $=0$ (20)
Since $\det(\mathbf{U}_{12}^{u\times u})=(-\lambda)^{u}$, and
$\det(\mathbf{U}_{12}^{-1})=\begin{vmatrix}-\frac{1}{\lambda}&0&\cdots&0\\ 0&-\frac{1}{\lambda}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&-\frac{1}{\lambda}\end{vmatrix}=\left(-\frac{1}{\lambda}\right)^{u}$ (21)
Given $\lambda\neq 0$, Eq. 20 is equivalent to
$\begin{vmatrix}\frac{u}{\lambda}-\lambda&\frac{u}{\lambda}&\cdots&\frac{u}{\lambda}\\ \frac{u}{\lambda}&\frac{u}{\lambda}-\lambda&\cdots&\frac{u}{\lambda}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{u}{\lambda}&\frac{u}{\lambda}&\cdots&\frac{u}{\lambda}-\lambda\end{vmatrix}^{v\times v}=0$ (22)
$\begin{vmatrix}u-\lambda^{2}&u&\cdots&u\\ u&u-\lambda^{2}&\cdots&u\\ \vdots&\vdots&\ddots&\vdots\\ u&u&\cdots&u-\lambda^{2}\end{vmatrix}^{v\times v}=0$ (23)
The roots of Eq. 23 are the eigenvalues of $u\cdot\mathbb{1}^{v\times v}$, which is a rank-1 matrix. The only non-zero eigenvalue of $u\cdot\mathbb{1}^{v\times v}$ is $uv$, leading to $\lambda^{2}=uv$. Thus, the two non-zero eigenvalues of $\mathbf{C}_{12}^{\prime}$ are ±$\pm\sqrt{uv}$, and the Pcor of $\mathbf{C}_{12}$ is
$\text{Pcor}(\mathbf{X}_{1},\mathbf{X}_{2})=\frac{1}{\sqrt{m}}\sqrt{\frac{2uv}{m-1}}$ (24)
Combining Eq. 15, 16, and 24, the overall inter-party correlation (Icor) is derived as
$\text{Icor}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=\frac{1}{2}\left({2*\text{Pcor}(\mathbf{X}_{1},\mathbf{X}_{2})-\text{Pcor}(\mathbf{X}_{1},\mathbf{X}_{1})-\text{Pcor}(\mathbf{X}_{2},\mathbf{X}_{2})}\right)$ (25)
$=\frac{1}{\sqrt{m}}\sqrt{\frac{2uv}{m-1}}-\frac{1}{\sqrt{m}}\sqrt{\frac{u^{2}+v^{2}}{m-1}}$
$=\frac{1}{\sqrt{m(m-1)}}\left(\sqrt{2uv}-\sqrt{u^{2}+v^{2}}\right)$
By the AM-HM inequality, $\text{Icor}(\mathbf{X}_{1},\mathbf{X}_{2})\leq 0$, with equality if and only if $u=v=\frac{m}{2}$.
For the lower bound, given $u=m-v$, expressing Icor in terms of $v$ gives
$\text{Icor}(\mathbf{X}_{1},\mathbf{X}_{2})=\frac{1}{\sqrt{m(m-1)}}\left(\sqrt{2mv-2v^{2}}-\sqrt{2v^{2}-2mv+m^{2}}\right)$ (26)
Letting $t=2mv-2v^{2}(v\in[0,m])$, it holds $t\in[0,m^{2}/2]$. Consider the function $f(t)=\sqrt{t}-\sqrt{m^{2}-t}$, the derivative of $f(t)$ ensures
$f^{\prime}(t)$ $=\frac{1}{2\sqrt{t}}+\frac{1}{2\sqrt{m^{2}-t}}$ (27)
$=\frac{\sqrt{m^{2}-t}-\sqrt{t}}{2\sqrt{t(m^{2}-t)}}\geq 0$
Therefore, $f(t)$ and $\text{Icor}(\mathbf{X}_{1},\mathbf{X}_{2})$ are monotonically non-descreasing w.r.t. $t$. The lower bound of Icor is reached when $t=0$, yielding
$\text{Icor}(\mathbf{X}_{1},\mathbf{X}_{2})\geq-\frac{m}{\sqrt{m(m-1)}}$ (28)
This condition holds if and only if $v=0$ or $v=m$.
Theorem 2.
[…]
@@
$\mathbb{E}[Z_{i}]=\sum_{j=1}^{m}\mathbb{E}[Y_{ij}]=\sum_{j=1}^{m}\phi_{j}(v)\mathbb{E}[r_{i}]=\mathbb{E}[r_{i}]\sum_{j=1}^{m}\phi(j)$ (30)
@@
$\mathbb{E}[Z_{i}]=\mathbb{E}[r_{i}]v(\mathcal{A})$ (31)
Since the property of Dirichlet distribution asserts that $\alpha_{i}\approx\mathbb{E}[r_{i}]$, it holds
$\mathbb{E}[Z_{i}]\propto\alpha_{i}v(\mathcal{A})$ (32)
@@
Suppose we have variables $X_{1},\dots,X_{K}$ following the Dirichlet distribution, denoted as $\text{Dir}({\alpha},\dots,{\alpha})$. Leveraging the inherent properties of the Dirichlet distribution, we can formulate the variance $\text{Var}(X_{i})$ for all $i\in[1,K]$ as
$\text{Var}(X_{i})=\frac{K-1}{K^{2}(K{\alpha}+1)}$ (34)
@@
$\mathbb{E}[\text{Var}(X_{i})]=\frac{K-1}{K^{2}(K{\alpha}+1)}=\sigma$ (35)
Recognizing that for a Dirichlet distribution $\sigma>0$ holds, we can transform the above equation to express ${\alpha}$ in terms of $\sigma$:
${\alpha}=\frac{K-1-K^{2}\sigma}{K^{3}\sigma}$ (36)
@@ Appendix B Details of VFL Algorithms
In this section, we provide a detailed comparison of the existing VFL algorithms as an extension of Section 4.1. This section critically reviews current VFL algorithms, with a focus on accuracy, efficiency, and communication size. VertiBench concentrates on standard supervised learning tasks such as classification and regression within synchronized parties, summarized in Table 1. Notably, this benchmark excludes studies exploring different VFL aspects such as privacy Jin et al. [2021], fairness Qi et al. [2022], data pricing Jiang et al. [2022], asynchronization Zhang et al. [2021b], Hu et al. [2019], Zhang et al. [2021a], latency Fu et al. [2022a], and other tasks like unsupervised learning Chang et al. [2020], matrix factorization Li et al. [2021b], multi-task learning Chen and Zhang [2022], and coreset construction Huang et al. [2022]. While most VFL algorithms presume accurate inter-party data linking, we adopt this approach in VertiBench, despite recent contrary findings Wu et al. [2022a], Nock et al. [2021] that this assumption may not be true.
The mainstream existing methods can be classified into two categories: ensemble-based and split-based. The distinguishing factor lies in the independent prediction capability of each party. Ensemble-based methods involve parties each maintaining a full model for local feature prediction, with collaborative ensemble methods during training, while split-based methods require each party to hold a partial model forming different inference stages of the full model. Consequently, split-based partial models cannot perform independent inference. For split-based models, our focus is on advanced models such as neural networks (NNs) and gradient boosting decision trees (GBDTs) Chen and Guestrin [2016], though VertiBench can accommodate various models Hardy et al. [2017], Gu et al. [2020]. Split-NN-based models are trained by transferring representations and gradients, while split-GBDT-models are trained by transferring gradients and histograms. While acknowledging the existence of various VFL algorithm subtypes Li et al. [2022b], our current study primarily addresses the major types of VFL algorithms. The exploration of other subtypes remains a subject for future research.
@@ B.2 Split-NN-based VFL Algorithms
FedOnce Wu et al. [2022b], a split-based VFL approach, distinguishes itself by its single-round communication protocol. Secondary parties initially engage in unsupervised learning, aiming to predict noise and subsequently develop local representations. These representations are then transferred to the primary party, which utilizes them in training a SplitNN model on the primary dataset. This method notably reduces both the communication size and the synchronization overhead.